OpenStudy (anonymous):

identify the solution(s) of the system of equations, if any 3x+2y=4 5x-4y=3

OpenStudy (mathmale):

Even though you may have several alternative answers from which to choose, it'd still make sense and be worth the effort to solve this system of linear equations yourself. Among the various methods of doing so are: 1. elimination by addition/subtraction 2. elimination by substitution 3. graphing 4. Gauss-Jordan approach Have you considered finding your own solution? If so, which approach would you consider to be most appropriate here?

OpenStudy (anonymous):

2y+3x-3x=4-3x 2y=-3x+4 y=-1.5x+0.5 plug in to other equation to find x then plug x into any equation to find y

OpenStudy (mathmale):

@yours925: Please become involved. Have you considered the questions I asked you a few minutes ago?

OpenStudy (anonymous):

multiplied by 2 we get 6x+4y=8 5x-4y-3 substracting we get x=1 &putting value we get y=1/2

OpenStudy (anonymous):

multiplied by 2 we get 6x+4y=8 5x-4y-3 substracting we get x=1 &putting value we get y=1/2

OpenStudy (anonymous):

helped?